# They See Me Rolling: High-Speed Event Vision-Based Tactile Roller Sensor for Large Surface Inspection

### v2

@@ I Introduction
Surface metrology and surface inspection are crucial elements in quality assurance across diverse industries, particularly aerospace and automotive manufacturing. Precise inspection is required to identify characteristics like paint quality, coating integrity, and subtle defects such as cracks, nicks, and dents [futterlieb2017air, yusra, Yuan2017]. Often, achieving a resolution of 0.1 mm or lower is necessary to accurately classify these features and ensure component integrity and safety [Agarwal2023]. Traditional contact-based methods, including high-precision profilometers [sekatskii2019analysis, Villarrubia1997] or microscopic techniques [Russell2001, Yang2018, Jonkman2020], offer high resolution locally but become exceedingly time-consuming when applied to large surface areas due to their sequential, point-by-point or small-patch measurement nature. Non-contact optical methods, such as cameras, laser scanners, or structured light systems [yusra, Salah2023, deng2024review, laser, zhang2018high, ramadan_vision-guided_2024], can significantly accelerate inspection by capturing data over wider areas. However, these methods often lack robustness; their performance can be compromised by variations in ambient lighting, motion blur when attempting high-speed scanning, or challenging surface optical properties like high reflectivity or transparency [Hausler2011]. This highlights a critical need for sensing modalities that combine high resolution and robustness with the capability for rapid inspection of large, potentially complex surfaces. Capturing the full 3D geometry is particularly important, as it enables the quantification of volumetric defects (e.g., dent depth and volume) and allows for direct comparison against a target CAD model, a critical step in modern quality assurance.
Vision-Based Tactile Sensors (VBTSs), particularly those derived from the GelSight concept [Yuan2017, Johnson2011, li2020review], have emerged as a powerful contact-based alternative offering a unique blend of high-resolution geometric and textural sensing. These sensors typically employ a camera to view the deformation of a soft, reflective elastomer membrane pressed against a target surface. This principle allows for detailed surface capture, often with inherent robustness to external lighting conditions due to internal illumination [Yuan2017, Dong2017]. Numerous variations of GelSight have been developed, enhancing capabilities for robotic manipulation, complex shape capture, and integration into compact form factors [Donlon2018, Liu2022, Zhao2023, Jing2025].
Despite their advantages for detailed local analysis, the standard planar configuration of most VBTSs necessitates a discrete ”press-and-lift” interaction mode. This process, where the sensor is repeatedly pressed, lifted, and moved, becomes prohibitively time-consuming when inspecting large surface areas, requiring numerous measurements to be stitched together [Agarwal2023], thereby limiting their use for efficient large-scale tactile measurement.
@@
Neuromorphic vision sensors, or event cameras, provide a promising technological solution to overcome the speed and motion blur limitations of conventional VBTSs [Gallego2022, Huang2022]. Unlike frame-based cameras, event cameras feature pixels that respond independently and asynchronously to logarithmic brightness changes, generating data with microsecond-level temporal resolution, high dynamic range, and inherent resistance to motion blur [Brandli2014, Posch2011, Lichtsteiner2008]. These characteristics are ideally suited for capturing the rapid, dynamic deformations occurring at the contact interface of a high-speed tactile sensor. Crucially, integrating event cameras into tactile sensing holds considerable potential for achieving high-resolution continuous tactile sensing at higher speeds with significantly lower computational power, particularly for surfaces with sparse details, compared to conventional VBTSs. While Neuromorphic VBTSs (NVBTS) have been explored for applications like robotic grasping and slip detection (e.g., Evetac [Funk2024]) or contact analysis [Ward-Cherrier2020, Sajwani2023, Faris2023, neutac_salah_2024], their potential for enabling high-speed, continuous 3D large surface reconstruction has not been realized.
To enable continuous tactile measurement, various designs moving beyond the static planar configuration have emerged. Attempts at direct sliding of planar sensors face significant friction, wear, and motion blur challenges [potdar_high-speed_2024]. Roller-based [Cao2023, Shimonomura2021] and belt-based [mirzaee_gelbelt_2025] VBTS designs replace the static contact with a dynamic rolling interface, allowing continuous data acquisition. TouchRoller, for example, demonstrated 2D texture mapping over moderate areas [Cao2023], while GelBelt provides a larger contact patch via its belt mechanism [mirzaee_gelbelt_2025]. Despite these mechanical advancements for measurement continuity, these designs face limitations. Reconstructing accurate 3D geometry from the depth-variant information gathered by cylindrical rollers is complex, primarily because the contact depth is not uniform across the curved sensing area [Li2023]. More critically, the operational speed of systems relying on conventional cameras is constrained. Frame rate limitations and motion blur become significant issues as scanning speed increases, particularly for VBTSs that require close-up imaging with a shallow depth of field to resolve fine surface details. Consequently, current roller/belt tactile sensors are restricted to relatively low velocities (e.g., typically well below 0.1 m/s for 3D reconstruction [Cao2023, mirzaee_gelbelt_2025]), insufficient for truly high-throughput industrial inspection of large parts. A significant gap exists for a tactile sensing system capable of combining the continuous scanning geometry of a roller with the high temporal fidelity and motion robustness of event-based vision for large-scale surface measurement.
@@ II Related Works
Reconstructing a surface’s three-dimensional geometry is crucial in many applications, from automated manufacturing to quality control. Mechanical profilometers have long been used for high-precision surface measurements, meticulously tracing the surface point-by-point. However, this inherent sequential nature makes them slow, especially when dealing with complex geometries or large areas [Villarrubia1997]. Non-contact optical techniques, like laser scanning and structured light, offer a significant speed advantage by rapidly capturing surface data over a wider area [Salvi2004, Sansoni2009]. These methods boast resolutions around 0.1 mm [Johnson2011], but their reliance on light makes them susceptible to issues with highly reflective or transparent materials, as well as variations in ambient lighting conditions [Hausler2011]. Furthermore, certain laser scanning systems, particularly higher-power ones used for speed or range, can introduce eye-safety concerns requiring adherence to safety standards and protocols [IEC60825].
For applications demanding the utmost precision, techniques like white light interferometry, confocal microscopy, and atomic force microscopy (AFM) are utilized [Russell2001, Yang2018, Jonkman2020]. These methods can resolve surface details down to the nanometer scale, revealing intricate textures and features. However, they often involve complex setups, lengthy measurement times, and significant expense, limiting their practical use for large-scale inspection tasks.
@@ II-B Vision-Based and Neuromorphic Tactile Sensing
Vision-Based Tactile Sensors (VBTSs) have emerged as compelling alternatives for surface sensing, offering robustness, versatility, and ease of use. The GelSight sensor pioneered this field, making use of a soft, deformable surface with camera-based observation to extract detailed surface information utilizing photometric stereo [Johnson2011]. By analyzing the deformation of the elastomer, typically coated with a reflective layer, VBTSs can capture high-resolution data on surface geometry and texture, even under varying lighting [Yuan2017, Dong2017]. From the original GelSight concept, a diverse range of VBTS designs have emerged to enhance performance and broaden application: GelSlim for more compact integrations [Donlon2018], GelSight Fin Ray for soft compliant grippers [Liu2022], and GelSight Svelte for human shape mimicking [Zhao2023], among others. Simulation tools have also evolved in parallel, accelerating the design and deployment of novel VBTS solutions [Agarwal2021, Wang2022, Akinola2025, Zhong2025].
@@
Despite these advantages, conventional VBTSs face challenges for continuous large-area scanning. The rigid backing of the elastomer limits its ability to slide smoothly across surfaces, leading to distortions and potential damage to the sensing membrane. As a result, most operate in a discrete “press-and-lift” mode, which is slow for large-area inspection; for example, the commercial GelSight Mobile™ Series 2 requires approximately 0.1 seconds of exposure time per measurement patch [GelSightMobile2024], making large-area coverage time-consuming [Agarwal2023]. Attempts at continuous sliding, such as with the DIGIT sensor for high-speed Braille reading [potdar_high-speed_2024], have shown potential but also revealed limitations: reliance on a conventional camera leads to motion blur at higher speeds, reducing recognition accuracy, while friction during sliding can cause rapid wear or damage to the elastomer. Protective layers can mitigate wear but may introduce optical diffusion, limiting fine texture resolution, and require precise contact control to ensure consistent data acquisition.
@@
Neuromorphic vision sensors, or event cameras, offer a transformative solution to these speed and motion blur limitations [Gallego2022]. Operating asynchronously at the pixel level, they detect brightness changes with microsecond temporal resolution, high dynamic range, low latency, and inherent suppression of motion blur [Brandli2014, Posch2011, Lichtsteiner2008]. These properties are well-suited to capturing the rapid, fine-scale elastomer deformations that occur during high-speed tactile scanning. Several Neuromorphic VBTSs (NVBTSs) have leveraged these advantages: Gaussian processes on event data for contact force estimation [Naeini2020], spiking neural networks for texture classification (NeuroTac) [Ward-Cherrier2020], event-based marker tracking for slip detection (Evetac) [Funk2024], graph neural networks for contact angle prediction (Tactigraph) [Sajwani2023], and event-based image representations for proprioception and contact analysis [Faris2023]. While these works validate the potential of NVBTSs, they have primarily focused on grasping-related tasks (slip, force, contact classification) or local texture analysis. Moreover, for these use cases, the algorithms used in combination with most NVBTS often result in decreased spatial resolution due to their reliance on tracking sparse markers, which is detrimental for surface reconstruction and depth estimation [Funk2024].
@@ II-C Roller-Based and Large Surface Tactile Sensors
Roller-based tactile sensors represent a promising avenue for achieving continuous tactile sensing, particularly for large surfaces where traditional planar sensors are inefficient. Early work explored cylindrical designs for specific detection tasks, such as the sensor developed by Shimonomura et al. [Shimonomura2021] for detecting foreign bodies like food impurities during rolling contact. Similarly, studies in robotic in-hand manipulation have successfully utilized a roller grasper, often equipped with high-resolution tactile feedback, for tasks like object insertion and regrasping [Lepert2023, Yuan2025], demonstrating the utility of rolling contact in dynamic scenarios beyond simple surface mapping.
@@
Transitioning towards surface property characterization, TouchRoller [Cao2023] demonstrated the potential for mapping surface textures by continuously rolling a sensor element across a fabric sample. It achieved coverage of an 8cm $\times$ 11cm area in 10 seconds, yielding a 2D texture map. Notably, its operational speed was limited to approximately 11 mm/s; the authors noted that higher speeds resulted in motion blur within the conventional camera frames, hindering performance. This motion blur problem is particularly acute for VBTSs, which require close-up imaging to resolve fine surface details. Operating at such close proximity means that even moderate rolling speeds translate into high velocities of features across the image sensor plane; during the camera’s finite exposure time, these features traverse a larger number of pixels, resulting in significant motion blur compared to imaging from a distance. Furthermore, TouchRoller did not perform 3D surface reconstruction. Addressing 3D reconstruction, GelRoller [zhang_gelroller_2024] employed a cylindrical design integrating photometric stereo techniques for local surface normal estimation from single frames. For reconstructing larger surfaces, they proposed stitching multiple local reconstructions using Iterative Closest Point (ICP) registration. While demonstrating 3D capability, this work did not explicitly analyze or optimize the operational speed of the rolling process.
@@
A fundamental challenge inherent to cylindrical roller designs, however, is that the tactile information gathered varies significantly with contact depth due to the sensor’s curvature. This leads to inconsistencies and potential distortions in the acquired tactile images, complicating accurate reconstruction. Researchers have explored methods to mitigate this depth inconsistency. Li et al. [Li2023] proposed a cyclic image fusion strategy, processing information from different contact phases within the rolling sequence using wavelet transforms to enhance information extraction and reduce bias from the curved contact.
@@
An alternative approach addresses the issue mechanically: the GelBelt sensor [mirzaee_gelbelt_2025] utilizes a flat elastomeric belt moving between two wheels, creating a larger, more uniform contact area akin to a conveyor. This design avoids the depth variation inherent in direct cylindrical contact and enabled continuous sensing at higher reported speeds, up to 45 mm/s. However, this design comes with some drawbacks. It relies on sliding motion to achieve high-speed operation, using an air gap and a smooth surface to reduce friction. While this allows GelBelt to move quickly, it creates challenges for lighting and can lead to wear over time, which may affect the system’s accuracy and repeatability. The design also struggles with curved surfaces. Its large, flat structure makes it hard to access concave areas, and it doesn’t perform well on convex surfaces either. In this case, the flat sensor doesn’t match the shape of the curved surface, leading to a limited sensing area, similar to problems seen in roller-based systems.
@@
Furthermore, despite these advances in measurement continuity, including GelBelt’s increased speed, systems relying on conventional cameras remain fundamentally constrained by the trade-off between speed and accuracy due to motion blur. As demonstrated by GelBelt’s own results showing decreased accuracy at higher speeds [mirzaee_gelbelt_2025], pushing scanning velocities further significantly degrades the quality of the captured tactile information and subsequent reconstructions. While naively increasing camera frame rates might seem like a straightforward solution to combat motion blur, it introduces substantial challenges regarding the increased computational complexity required for real-time processing, significantly larger data storage requirements, and higher energy consumption, potentially limiting practical deployment. In turn, the potential for using event-based vision in roller sensors to achieve similar resolution at low computational power and significantly higher speeds for continuous tactile sensing is considerable. However, this capability has yet to be demonstrated, particularly for demanding large-scale inspection tasks.

### v3

@@ I Introduction
Surface metrology and surface inspection are crucial elements in quality assurance across diverse industries, particularly aerospace and automotive manufacturing. Precise inspection is required to identify characteristics like paint quality, coating integrity, and subtle defects such as cracks, nicks, and dents [18, 1, 60]. Often, achieving a resolution of 0.1 mm or lower is necessary to accurately classify these features and ensure component integrity and safety [2]. Traditional contact-based methods, including high-precision profilometers [48, 52] or microscopic techniques [41, 57, 27], offer high resolution locally but become exceedingly time-consuming when applied to large surface areas due to their sequential, point-by-point or small-patch measurement nature. Non-contact optical methods, such as cameras, laser scanners, or structured light systems [1, 44, 12, 22, 61, 39], can significantly accelerate inspection by capturing data over wider areas. However, these methods often lack robustness; their performance can be compromised by variations in ambient lighting, motion blur when attempting high-speed scanning, or challenging surface optical properties like high reflectivity or transparency [23]. This highlights a critical need for sensing modalities that combine high resolution and robustness with the capability for rapid inspection of large, potentially complex surfaces. Capturing the full 3D geometry is particularly important, as it enables the quantification of volumetric defects (e.g., dent depth and volume) and allows for direct comparison against a target CAD model, a critical step in modern quality assurance.
Vision-Based Tactile Sensors (VBTSs), particularly those derived from the GelSight concept [60, 26, 30], have emerged as a powerful contact-based alternative offering a unique blend of high-resolution geometric and textural sensing. These sensors typically employ a camera to view the deformation of a soft, reflective elastomer membrane pressed against a target surface. This principle allows for detailed surface capture, often with inherent robustness to external lighting conditions due to internal illumination [60, 13]. Numerous variations of GelSight have been developed, enhancing capabilities for robotic manipulation, complex shape capture, and integration into compact form factors [14, 34, 63, 55].
Despite their advantages for detailed local analysis, the standard planar configuration of most VBTSs necessitates a discrete ”press-and-lift” interaction mode. This process, where the sensor is repeatedly pressed, lifted, and moved, becomes prohibitively time-consuming when inspecting large surface areas, requiring numerous measurements to be stitched together [2], thereby limiting their use for efficient large-scale tactile measurement.
@@
Neuromorphic vision sensors, or event cameras, provide a promising technological solution to overcome the speed and motion blur limitations of conventional VBTSs [19, 25]. Unlike frame-based cameras, event cameras feature pixels that respond independently and asynchronously to logarithmic brightness changes, generating data with microsecond-level temporal resolution, high dynamic range, and inherent resistance to motion blur [9, 37, 33]. These characteristics are ideally suited for capturing the rapid, dynamic deformations occurring at the contact interface of a high-speed tactile sensor. Crucially, integrating event cameras into tactile sensing holds considerable potential for achieving high-resolution continuous tactile sensing at higher speeds with significantly lower computational power, particularly for surfaces with sparse details, compared to conventional VBTSs. While Neuromorphic VBTSs (NVBTS) have been explored for applications like robotic grasping and slip detection (e.g., Evetac [17]) or contact analysis [54, 42, 15, 45], their potential for enabling high-speed, continuous 3D large surface reconstruction has not been realized.
To enable continuous tactile measurement, various designs moving beyond the static planar configuration have emerged. Attempts at direct sliding of planar sensors face significant friction, wear, and motion blur challenges [38]. Roller-based [10, 50] and belt-based [35] VBTS designs replace the static contact with a dynamic rolling interface, allowing continuous data acquisition. TouchRoller, for example, demonstrated 2D texture mapping over moderate areas [10], while GelBelt provides a larger contact patch via its belt mechanism [35]. Despite these mechanical advancements for measurement continuity, these designs face limitations. Reconstructing accurate 3D geometry from the depth-variant information gathered by cylindrical rollers is complex, primarily because the contact depth is not uniform across the curved sensing area [32]. More critically, the operational speed of systems relying on conventional cameras is constrained. Frame rate limitations and motion blur become significant issues as scanning speed increases, particularly for VBTSs that require close-up imaging with a shallow depth of field to resolve fine surface details. Consequently, current roller/belt tactile sensors are restricted to relatively low velocities (e.g., typically well below 0.1 m/s for 3D reconstruction [10, 35]), insufficient for truly high-throughput industrial inspection of large parts. A significant gap exists for a tactile sensing system capable of combining the continuous scanning geometry of a roller with the high temporal fidelity and motion robustness of event-based vision for large-scale surface measurement.
@@ II Related Works
Reconstructing a surface’s three-dimensional geometry is crucial in many applications, from automated manufacturing to quality control. Mechanical profilometers have long been used for high-precision surface measurements, meticulously tracing the surface point-by-point. However, this inherent sequential nature makes them slow, especially when dealing with complex geometries or large areas [52]. Non-contact optical techniques, like laser scanning and structured light, offer a significant speed advantage by rapidly capturing surface data over a wider area [46, 47]. These methods boast resolutions around 0.1 mm [26], but their reliance on light makes them susceptible to issues with highly reflective or transparent materials, as well as variations in ambient lighting conditions [23]. Furthermore, certain laser scanning systems, particularly higher-power ones used for speed or range, can introduce eye-safety concerns requiring adherence to safety standards and protocols [28].
For applications demanding the utmost precision, techniques like white light interferometry, confocal microscopy, and atomic force microscopy (AFM) are utilized [41, 57, 27]. These methods can resolve surface details down to the nanometer scale, revealing intricate textures and features. However, they often involve complex setups, lengthy measurement times, and significant expense, limiting their practical use for large-scale inspection tasks.
@@ II-B Vision-Based and Neuromorphic Tactile Sensing
Vision-Based Tactile Sensors (VBTSs) have emerged as compelling alternatives for surface sensing, offering robustness, versatility, and ease of use. The GelSight sensor pioneered this field, making use of a soft, deformable surface with camera-based observation to extract detailed surface information utilizing photometric stereo [26]. By analyzing the deformation of the elastomer, typically coated with a reflective layer, VBTSs can capture high-resolution data on surface geometry and texture, even under varying lighting [60, 13]. From the original GelSight concept, a diverse range of VBTS designs have emerged to enhance performance and broaden application: GelSlim for more compact integrations [14], GelSight Fin Ray for soft compliant grippers [34], and GelSight Svelte for human shape mimicking [63], among others. Simulation tools have also evolved in parallel, accelerating the design and deployment of novel VBTS solutions [3, 53, 4, 64].
@@
Despite these advantages, conventional VBTSs face challenges for continuous large-area scanning. The rigid backing of the elastomer limits its ability to slide smoothly across surfaces, leading to distortions and potential damage to the sensing membrane. As a result, most operate in a discrete “press-and-lift” mode, which is slow for large-area inspection; for example, the commercial GelSight Mobile™ Series 2 requires approximately 0.1 seconds of exposure time per measurement patch [21], making large-area coverage time-consuming [2]. Attempts at continuous sliding, such as with the DIGIT sensor for high-speed Braille reading [38], have shown potential but also revealed limitations: reliance on a conventional camera leads to motion blur at higher speeds, reducing recognition accuracy, while friction during sliding can cause rapid wear or damage to the elastomer. Protective layers can mitigate wear but may introduce optical diffusion, limiting fine texture resolution, and require precise contact control to ensure consistent data acquisition.
@@
Neuromorphic vision sensors, or event cameras, offer a transformative solution to these speed and motion blur limitations [19]. Operating asynchronously at the pixel level, they detect brightness changes with microsecond temporal resolution, high dynamic range, low latency, and inherent suppression of motion blur [9, 37, 33]. These properties are well-suited to capturing the rapid, fine-scale elastomer deformations that occur during high-speed tactile scanning. Several Neuromorphic VBTSs (NVBTSs) have leveraged these advantages: Gaussian processes on event data for contact force estimation [6], spiking neural networks for texture classification (NeuroTac) [54], event-based marker tracking for slip detection (Evetac) [17], graph neural networks for contact angle prediction (Tactigraph) [42], spiking neural networks for robust Braille reading in noisy environments [56], contact status and force estimation (GelEvent [58]), and event-based image representations for proprioception and contact analysis [15]. While these works validate the potential of NVBTSs, they have primarily focused on grasping-related tasks (slip, force, contact classification) or local texture analysis. Moreover, for these use cases, the algorithms used in combination with most NVBTS often result in decreased spatial resolution due to their reliance on tracking sparse markers, which is detrimental for surface reconstruction and depth estimation [17].
@@ II-C Roller-Based and Large Surface Tactile Sensors
Roller-based tactile sensors represent a promising avenue for achieving continuous tactile sensing, particularly for large surfaces where traditional planar sensors are inefficient. Early work explored cylindrical designs for specific detection tasks, such as the sensor developed by Shimonomura et al. [50] for detecting foreign bodies like food impurities during rolling contact. Similarly, studies in robotic in-hand manipulation have successfully utilized a roller grasper, often equipped with high-resolution tactile feedback, for tasks like object insertion and regrasping [29, 59], demonstrating the utility of rolling contact in dynamic scenarios beyond simple surface mapping.
@@
Transitioning towards surface property characterization, TouchRoller [10] demonstrated the potential for mapping surface textures by continuously rolling a sensor element across a fabric sample. It achieved coverage of an 8cm $\times$ 11cm area in 10 seconds, yielding a 2D texture map. Notably, its operational speed was limited to approximately 11 mm/s; the authors noted that higher speeds resulted in motion blur within the conventional camera frames, hindering performance. This motion blur problem is particularly acute for VBTSs, which require close-up imaging to resolve fine surface details. Operating at such close proximity means that even moderate rolling speeds translate into high velocities of features across the image sensor plane; during the camera’s finite exposure time, these features traverse a larger number of pixels, resulting in significant motion blur compared to imaging from a distance. Furthermore, TouchRoller did not perform 3D surface reconstruction. Addressing 3D reconstruction, GelRoller [62] employed a cylindrical design integrating photometric stereo techniques for local surface normal estimation from single frames. For reconstructing larger surfaces, they proposed stitching multiple local reconstructions using Iterative Closest Point (ICP) registration. While demonstrating 3D capability, this work did not explicitly analyze or optimize the operational speed of the rolling process.
@@
A fundamental challenge inherent to cylindrical roller designs, however, is that the tactile information gathered varies significantly with contact depth due to the sensor’s curvature. This leads to inconsistencies and potential distortions in the acquired tactile images, complicating accurate reconstruction. Researchers have explored methods to mitigate this depth inconsistency. Li et al. [32] proposed a cyclic image fusion strategy, processing information from different contact phases within the rolling sequence using wavelet transforms to enhance information extraction and reduce bias from the curved contact.
@@
An alternative approach addresses the issue mechanically: the GelBelt sensor [35] utilizes a flat elastomeric belt moving between two wheels, creating a larger, more uniform contact area akin to a conveyor. This design avoids the depth variation inherent in direct cylindrical contact and enabled continuous sensing at higher reported speeds, up to 45 mm/s. Building on this hardware, GelSLAM [24] introduced a real-time tactile SLAM system capable of estimating object pose and reconstructing global shape by fusing local surface normals. However, this design comes with some drawbacks. It relies on sliding motion to achieve high-speed operation, using an air gap and a smooth surface to reduce friction. While this allows GelBelt to move quickly, it creates challenges for lighting and can lead to wear over time, which may affect the system’s accuracy and repeatability. The design also struggles with curved surfaces. Its large, flat structure makes it hard to access concave areas, and it doesn’t perform well on convex surfaces either. In this case, the flat sensor doesn’t match the shape of the curved surface, leading to a limited sensing area, similar to problems seen in roller-based systems.
@@
Furthermore, despite these advances in measurement continuity, including GelBelt’s increased speed, systems relying on conventional cameras remain fundamentally constrained by the trade-off between speed and accuracy due to motion blur. As demonstrated by GelBelt’s own results showing decreased accuracy at higher speeds [35], pushing scanning velocities further significantly degrades the quality of the captured tactile information and subsequent reconstructions. While naively increasing camera frame rates might seem like a straightforward solution to combat motion blur, it introduces substantial challenges regarding the increased computational complexity required for real-time processing, significantly larger data storage requirements, and higher energy consumption, potentially limiting practical deployment. In turn, the potential for using event-based vision in roller sensors to achieve similar resolution at low computational power and significantly higher speeds for continuous tactile sensing is considerable. However, this capability has yet to be demonstrated, particularly for demanding large-scale inspection tasks.